OpenStudy (anonymous):

find th eformula for the following sequence: 3,15,75,...

OpenStudy (anonymous):

I know it is 3*5=15, then 15*5=75, then 75*5=375. But I can't get the formula, let me think a little bit more

OpenStudy (anonymous):

a0=3 a1 = 5*a0 a2 = 5*5*a1 = 3(5^2) a3 = (5^3) a0 etc.